# Tac-Man: Tactile-Informed Prior-Free Manipulation of Articulated Objects

### v1

@@ Abstract
Integrating robotics into human-centric environments such as homes, necessitates advanced manipulation skills as robotic devices will need to engage with articulated objects like doors and drawers. Key challenges in robotic manipulation are the unpredictability and diversity of these objects’ internal structures, which render models based on priors, both explicit and implicit, inadequate. Their reliability is significantly diminished by pre-interaction ambiguities, imperfect structural parameters, encounters with unknown objects, and unforeseen disturbances. Here, we present a prior-free strategy, Tac-Man, focusing on maintaining stable robot-object contact during manipulation. Utilizing tactile feedback, but independent of object priors, Tac-Man enables robots to proficiently handle a variety of articulated objects, including those with complex joints, even when influenced by unexpected disturbances. Demonstrated in both real-world experiments and extensive simulations, it consistently achieves near-perfect success in dynamic and varied settings, outperforming existing methods. Our results indicate that tactile sensing alone suffices for managing diverse articulated objects, offering greater robustness and generalization than prior-based approaches. This underscores the importance of detailed contact modeling in complex manipulation tasks, especially with articulated objects. Advancements in tactile sensors significantly expand the scope of robotic applications in human-centric environments, particularly where accurate models are difficult to obtain.
@@ I Introduction
Articulated objects, defined by interconnected parts that move in conjunction with each other, include a vast array of everyday items, such as doors, drawers, and appliances, as illustrated in LABEL:fig:teaser. Robots’ capabilities in the manipulation of various articulated objects are essential for effective integration into human-centric environments, assisting humans in diverse tasks while representing a significant challenge within robotics [1, 2].
To interact efficiently with these objects, robots traditionally rely on prior knowledge of their kinematics. This knowledge is often embedded either explicitly, as in kinematic models used for designing manipulation strategies via control or planning techniques [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20], or implicitly, within policy models developed through machine learning techniques [21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38]. However, acquiring precise prior knowledge about such objects presents several spatial and temporal challenges:
  • Ambiguity: Objects with similar external appearances may have vastly different internal structures, complicating the unique determination of their kinematics through perception. This issue is extensively explored in Zhu et al. [39].
@@
  • Obsolescence: Kinematic models can become outdated due to active or passive changes over time, leading to a gap in knowledge when executing manipulation tasks.
@@
Drawing inspiration from human tactile interaction, we propose a novel prior-free approach, Tac-Man. Tac-Man is centered around stable robot-object contact during manipulation, achieved through high-resolution tactile sensing. By integrating GelSight-inspired sensors [40, 41] into the robot’s gripper, Tac-Man adapts according to the tactile feedback during the interaction, allowing for dynamic adjustment during object handling. Without any prior-based modeling, Tac-Man addresses the challenges posed by ambiguity, imperfectness, unknowns, and obsolescence in object kinematics, which are prevalent in dynamic human-centric environments.
@@ II-B Manipulation with Implicit Priors
Imitation learning enables robots to acquire implicit priors through the observation of human demonstrations [21, 22, 23, 24, 25, 26, 27, 28, 29, 30], sometimes facilitated by tele-operational systems [42]. One major challenge in this approach is the collection of diverse and rich data from humans [43]. Although transfer methods [44, 45, 46, 47] have improved the applicability to similar objects, extending these priors to a wider range of objects remains a substantial hurdle.
@@ II-C Robot-Object Contact Modeling
II-D Tactile Sensors
Compared to 6D force/torque sensors commonly mounted at robots’ wrists, tactile sensors play a crucial role in directly measuring the state of robot-object contact [50, 51]. Commercially available solutions like the BioTac sensor [52] are known for their high sensitivity, though they often lack spatial resolution. Alternative technologies, such as force-sensitive resistors, piezoresistive materials, and strain gauges, offer flexibility in mounting and are useful for single-point force sensing [53, 54]. However, they do not provide detailed spatial information about contact. E-skins, recognized for their adaptability and multi-modal sensing capabilities, deliver high resolution and are a promising development [55, 56, 57, 58, 59, 60]. Yet, they are prone to manufacturing complexities, durability issues, and higher costs.
GelSight-type sensors, on the other hand, present a compelling alternative [61, 62, 63, 64, 65, 66, 67, 68, 69]. Characterized by pixel-level spatial resolution and cost-effective production, these sensors represent significant progress in tactile sensing technology. In our research, we have developed two GelSight-inspired tactile sensors, each uniquely mounted on the Robotiq 2F-85 gripper’s opposing sides. These sensors are custom-designed with distributed markers [41], enabling efficient feature correspondence between frames. This design improves the gripper’s ability to perceive fine details of contact, facilitating more nuanced manipulation of articulated objects.
@@ III-A Notation and Preliminaries
  • Bold small letters represent vectors (e.g., $\mathbf{v}$), while capital letters denote matrices (e.g., $M$). The subscript $[i]$, where $i=1,\dots,n$, on a bold small letter, refers to the $i$-th element of the vector. $I_{n\times n}$ is used to represent the identity matrix of dimension $n$.
  • The symbol $\hat{\cdot}$ indicates the axis of a frame. The generalized pose of a frame $\{i\}$ relative to another frame $\{j\}$ is described by a position vector $\mathbf{p}_{i}^{j}\in\mathbb{R}^{3}$ and a rotation matrix $R_{i}^{j}\in SO(3)$. For simplicity, when frame $\{j\}$ is equivalent to the world frame $\{w\}$, the superscript is omitted. The homogeneous transformation matrix $T_{i}^{j}$, integrating the position vector and rotation matrix, is given by:
    $T_{i}^{j}=\begin{bmatrix}R_{i}^{j}&\mathbf{p}_{i}^{j}\\ \begin{matrix}0&0&0\end{matrix}&1\end{bmatrix}.$
  • $\mathcal{K}_{ij}$ refers to a set of corresponding points from point clouds $C_{i}$ and $C_{j}$, formulated as:
    $\mathcal{K}_{ij}=\{(\mathbf{u},\mathbf{v})\mid\mathbf{u}\in C_{i},\mathbf{v}\in C_{j}\}.$
    The notation #\#$\cdot$# ⋅ is used to denote the number of elements in a set.
@@ III-B Problem Formulation
In this subsection, we formulate the process of tactile-informed prior-free manipulation of articulated objects. Consider an articulated object with a movable part, which is assumed to be graspable through a rigidly attached “handle.” We postulate that the object’s articulation restricts the part and the handle to a 1- Degree-of-freedom (DoF) trajectory. Let us denote the handle’s trajectory as $\mathcal{J}$, beginning at point $\mathcal{J}_{s}$ and concluding at $\mathcal{J}_{e}$. Initially, in the absence of priors, only $\mathcal{J}_{s}$ is known. A robotic gripper, positioned in a global pose $T_{1}$ is grasping the handle. The objective is to navigate the gripper through a series of poses, $T_{1},T_{2},T_{3},\dots$, to maneuver the handle towards $\mathcal{J}_{e}$. At any given pose $T_{i}$, the contact state $C_{i}$ between the object handle and the robotic gripper can be ascertained via a mapping function $f_{c}$:
$C_{i}=f_{c}(\mathcal{J},T_{i}).$ (1)
In practical terms, $C_{i}$ is directly measurable through tactile sensors. Importantly, the state prior to establishing any contact is denoted as $C_{0}$.
@@
To ensure successful manipulation, it is crucial that the contacts $C_{1:n}$ meet certain constraints, guaranteeing that the gripper can consistently grasp the handle throughout the process. In this paper, we characterize a state that meets these constraints as maintaining stable contact.
Stable contact has two features. First, it must ensure that the gripper maintains an intact grip on the handle throughout the manipulation process and prevents slippage. This concept is quantified by referencing the initial contact state $C_{0}$ and assessing subsequent contact states $C_{i}$. We employ the function $f_{d}(C_{i},C_{0})$ to calculate the deformation of $C_{i}$; this deformation must stay within the elastic deformation limits $D$ of the gripper material. Additionally, the shear deformation, as determined by the function $f_{s}(C_{i},C_{0})$, should not surpass $S$, a threshold dictated by the friction ratio between the handle and gripper material. The mathematical expression, therefore, is formulated as:
$\begin{bmatrix}f_{d}(C_{i},C_{0})\\ f_{s}(C_{i},C_{0})\end{bmatrix}\leq\begin{bmatrix}D\\ S\end{bmatrix}.$ (2)
Furthermore, a stable contact necessitates the gripper’s engagement with the object’s handle to ensure synchronized movement with the robot’s actions. To establish a reference for stable contact $C_{1}$, an initial predefined pose $T_{1}$ is set. This practice is common in the manipulation of articulated objects [8]. A contact $C_{i}$ is considered stable also requires its deviation from $C_{1}$ stays within a specified threshold $E$:
$f_{e}(C_{i},C_{1})\leq E,$ (3)
where $f_{e}(C_{i},C_{1})$ computes the difference between $C_{i}$ and $C_{1}$.
@@
Upon establishing stable contact, the manipulation process commences with a preliminary direction relative to frame $\{1\}$ for operating the articulated object. This direction, represented as $\mathcal{D}$, is encapsulated by a rotation matrix $R^{1}$ in $SO(3)$, detailed in Sec. VI-B. The homogeneous transformation matrix $T_{\mathcal{D}}$, resulting from moving along $\mathcal{D}$ by a distance $y$, is defined as:
$T_{\mathcal{D}}=\begin{bmatrix}I_{3\times 3}&R^{1}\mathbf{t}\\ \begin{matrix}0&0&0\end{matrix}&1\end{bmatrix},$ (4)
where $\mathbf{t}$, the translation vector, corresponds to the manipulation length along the negative y-axis as:
$\mathbf{t}=\begin{bmatrix}0\\ -y\\ 0\end{bmatrix}.$ (5)
The robot follows the preliminary direction, executing a series of gripper frame poses $T_{1:n}$ to guide the handle’s transition from $\mathcal{J}_{s}$ to $\mathcal{J}_{e}$ while ensuring continuous stable contact. For each incremental step $i=1,\dots,n-1$, the next pose $T_{i+1}$ is computed as follows:
$T_{i+1}=T_{i}T_{i+1}^{i}=T_{i}\underbracket{T_{\mathcal{D}}}_{\text{execution}}\underbracket{\operatorname*{arg\,min}_{T_{R_{i}}\in\mathbb{R}^{4\times 4}}\left\|f_{c}(T_{i}T_{R_{i}},\mathcal{J})-C_{1}\right\|_{2}}_{\text{recovery}}.$ (6)
@@
The execution term in Eq. 6 involves proceeding to move along the preliminary direction as described in Eq. 4. This navigation progresses under the assumption of no prior knowledge about the object handle’s trajectory, posing a risk of exceeding the tolerance limits necessary for maintaining stable contact. To balance this risk and maintain efficiency, it is crucial to carefully select the magnitude of $y$ in Eq. 4. The objective is to maximize $y$ while ensuring that the constraints for stable contact are met. This optimization can be mathematically represented as:
@@
subject to $f_{d}(f_{c}(T_{i}T_{\mathcal{D}},\mathcal{J}),C_{0})\leq D,$
$f_{s}(f_{c}(T_{i}T_{\mathcal{D}},\mathcal{J}),C_{0})\leq S,$
$f_{e}(f_{c}(T_{i}T_{\mathcal{D}},\mathcal{J}),C_{1})\leq E.$
@@
The recovery term in Eq. 6 implies that at each step, the robot adjusts its pose to minimize any deviation from the stable contact state $C_{1}$, which may arise due to the execution step.

### v3

@@ Abstract
Integrating robots into human-centric environments such as homes, necessitates advanced manipulation skills as robotic devices will need to engage with articulated objects like doors and drawers. Key challenges in robotic manipulation of articulated objects are the unpredictability and diversity of these objects’ internal structures, which render models based on object kinematics priors, both explicit and implicit, inadequate. Their reliability is significantly diminished by pre-interaction ambiguities, imperfect structural parameters, encounters with unknown objects, and unforeseen disturbances. Here, we present a prior-free strategy, Tac-Man, focusing on maintaining stable robot-object contact during manipulation. Without relying on object priors, Tac-Man leverages tactile feedback to enable robots to proficiently handle a variety of articulated objects, including those with complex joints, even when influenced by unexpected disturbances. Demonstrated in both real-world experiments and extensive simulations, it consistently achieves near-perfect success in dynamic and varied settings, outperforming existing methods. Our results indicate that tactile sensing alone suffices for managing diverse articulated objects, offering greater robustness and generalization than prior-based approaches. This underscores the importance of detailed contact modeling in complex manipulation tasks, especially with articulated objects. Advancements in tactile-informed approaches significantly expand the scope of robotic applications in human-centric environments, particularly where accurate models are difficult to obtain. See additional material at https://tacman-aom.github.io.
@@ I Introduction
Articulated objects, defined by interconnected parts that move in conjunction with each other, include a vast array of everyday items, such as doors, drawers, and appliances [1], as illustrated in LABEL:fig:teaser. A robot’s capability to manipulate articulated objects is essential for effective integration into human-centric environments, assisting humans in diverse tasks while representing a significant challenge within robotics [2, 3].
To interact efficiently with these objects, robots traditionally rely on prior knowledge of their kinematics. This knowledge is often embedded either explicitly, as in kinematic models used for designing manipulation strategies via control or planning techniques [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21], or implicitly, within policy models developed through machine learning techniques [22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38]. However, acquiring precise prior knowledge about such objects presents several spatial and temporal challenges:
  • Ambiguity: Objects with similar external appearances may have vastly different internal structures, complicating the unique determination of their kinematics through perception. This issue is extensively explored by Zhu et al. [39].
@@
  • Obsolescence: Kinematic models can become obsolescent due to active or passive changes over time, leading to a gap in knowledge when executing manipulation tasks.
@@
Drawing inspiration from human tactile interaction, we propose a novel tactile-informed prior-free manipulation approach, Tac-Man. Tac-Man is centered around stable robot-object contact during manipulation. By accurately sensing contact geometry information, Tac-Man leverages deviations from the initial state to inform the robot of both the direction and magnitude of adjustments needed to effectively manipulate the articulated object. Without any prior-based modeling, Tac-Man addresses the challenges posed by ambiguity, imperfectness, unknowns, and obsolescence in object kinematics, which are prevalent in dynamic human-centric environments.
@@ II-B Manipulation with Implicit Priors
Imitation learning enables robots to acquire implicit priors through the observation of human demonstrations [22, 23, 24, 25, 26, 27, 28, 29, 30], sometimes facilitated by tele-operational systems [41]. One major challenge in this approach is the collection of diverse and rich data from humans [42]. Although transfer methods [43, 44, 45] have improved the applicability to similar objects, extending these priors to a wider range of objects remains a substantial hurdle.
@@ II-C Robot-Object Contact Modeling
II-D Contact Sensing
Sensing of robot-object contacts is crucial for contact regulation [48, 49, 50, 51, 52, 53]. Decades of development have yielded various sensing technologies regarding this issue, broadly categorized as intrinsic or extrinsic styles. Intrinsic tactile sensors, such as 6D force/torque sensors [54], are integrated within the robot’s structure, providing indirect contact measurements through internal force sensing. Extrinsic tactile sensors, attached to the robot’s interacting surface, often measure contact information in a more direct manner, including conductive-fluid-based (e.g., BioTac [55]), resistor-based [56, 57], piezoelectricity-based [58, 59], capacitance-based [60, 61], and vision-based [62, 63, 64, 65] sensors, each with unique principles for measuring robot-object contacts.
The effectiveness of Tac-Man relies on timely compensation for contact deviations, which are modeled geometrically as the changes in the contact point positions, effectively informing both the direction and the magnitude of the necessary adjustments. Under this formulation, sensors with a primary focus on sensing contact geometry, such as GelSight [62], TacTip [63], and others, can be directly adopted within the Tac-Man framework. Additionally, sensors primarily designed to measure contact forces can also be utilized within our Tac-Man framework, provided that they are properly calibrated according to Hooke’s law [66] to obtain geometric information. A detailed discussion on sensor selection is available in Sec. VI.
In this research, we adopt the GelSight-type sensor for the pipeline illustration and experiment verification of the Tac-Man among the set of sensors with geometric measurements as one economical option. However, it should be noted that there are no methodological preferences among these sensors with geometric measurement ability, and the formulations should be very similar to each other, with only slight differences in the pre-processing procedure where contact deviations are extracted from the raw sensor data.
II-E Tactile Servoing
Tactile servoing, which involves adjusting robot poses in response to tactile signals, has been a significant area of research since its inception by Weiss et al. [67] and first implementation by Berger and Khosla [68]. It has evolved into two main approaches: model-based and learning-based. Model-based tactile servoing uses explicit models such as the inverse tactile model [69], the tactile Jacobian model [70, 71], the projector matrix [72, 73], or other tactile-based controllers [74, 75] to map tactile feedback onto robot actions. Learning-based tactile servoing leverages modern machine learning techniques, including imitation learning [76] and deep learning [77, 78, 53], to enable implicit mapping for complex robot manipulations. Tac-Man builds upon the fundamental principles of model-based tactile servoing, demonstrating its efficacy in the intricate manipulation of articulated objects.
@@ III-A Notation and Preliminaries
  • Bold lowercase letters represent vectors (e.g., $\boldsymbol{p}$), with the subscript $i$ denoting the $i$-th entry of the vector (e.g., $p_{i}$).
  • Bold uppercase letters represent matrices (e.g., $\boldsymbol{M}$). The identity matrix of dimension $n$ is denoted by $\boldsymbol{I}_{n\times n}$. The superscript $\cdot^{\mathsf{T}}$ indicates the transpose.
  • The symbol $\hat{\cdot}$ indicates the axis of a frame. The generalized pose of a frame $\{i\}$ relative to another frame $\{j\}$ is described by a position vector $\boldsymbol{p}_{i}^{j}\in\mathbb{R}^{3}$ and a rotation matrix $\boldsymbol{R}_{i}^{j}\in SO(3)$. For simplicity, when frame $\{j\}$ is equivalent to the world frame $\{w\}$, the superscript is omitted. The homogeneous transformation matrix $\boldsymbol{T}_{i}^{j}\in SE(3)$, integrating the position vector and rotation matrix, is given by:
    $\boldsymbol{T}_{i}^{j}=\begin{bmatrix}\boldsymbol{R}_{i}^{j}&\boldsymbol{p}_{i}^{j}\\ \begin{matrix}0&0&0\end{matrix}&1\end{bmatrix}.$
[…]
    $\mathcal{K}_{ij}=\{(\boldsymbol{u},\boldsymbol{v})\mid\boldsymbol{u}\in\mathcal{C}_{i},\boldsymbol{v}\in\mathcal{C}_{j}\}.$
@@ III-B Problem Formulation
In this subsection, we formulate the process of tactile-informed prior-free manipulation of articulated objects. Consider an articulated object with a movable part, which is assumed to be graspable through a rigidly attached “handle.” We postulate that the object’s articulation restricts the handle to a 1- Degree-of-freedom (DoF) trajectory. Let us denote the handle’s trajectory as $J$, beginning at point $J_{s}$ and concluding at $J_{e}$. Initially, in the absence of priors, only $J_{s}$ is known. A robotic gripper, positioned in a global pose $\boldsymbol{T}_{1}$ is grasping the handle. The objective is to navigate the gripper through a series of poses, $\boldsymbol{T}_{1},\boldsymbol{T}_{2},\boldsymbol{T}_{3},\dots$, to maneuver the handle towards $J_{e}$. At any given pose $\boldsymbol{T}_{i}$, the contact state $\mathcal{C}_{i}$ between the object handle and the robotic gripper can be ascertained via a mapping function $f_{c}$:
$\mathcal{C}_{i}=f_{c}(J,\boldsymbol{T}_{i}).$ (1)
In practical terms, $\mathcal{C}_{i}$ is directly measurable through tactile sensors. Importantly, the state prior to establishing any contact is denoted as $\mathcal{C}_{0}$.
@@
To ensure successful manipulation, it is crucial that the contacts $\mathcal{C}_{1:n}$ meet certain constraints, guaranteeing that the gripper can consistently grasp the handle throughout the process. In this paper, we characterize a state that meets these constraints as maintaining stable contact.
Stable contact is characterized by two essential features. First, the gripper has to maintain a secure grip on the handle without causing damage to itself, while also preventing slippage. This ensures that any relative pose changes between the grip and handle are accurately mirrored in the contact changes. This concept is quantified by referencing state prior to establishing any contact $\mathcal{C}_{0}$ and evaluating subsequent contact states $\mathcal{C}_{i}$. We use the function $f_{d}(\mathcal{C}_{i},\mathcal{C}_{0})$ to calculate the deformation of $\mathcal{C}_{i}$; this deformation must remain within the elastic limits $\boldsymbol{e}$ of the gripper material. Moreover, the shear deformation, as calculated by the function $f_{s}(\mathcal{C}_{i},\mathcal{C}_{0})$, should not exceed $s$, a threshold determined by the friction ratio between the handle and gripper material. Formally, this can be expressed as:
$\begin{bmatrix}f_{e}(\mathcal{C}_{i},\mathcal{C}_{0})\\ f_{s}(\mathcal{C}_{i},\mathcal{C}_{0})\end{bmatrix}\leq\begin{bmatrix}\boldsymbol{e}\\ s\end{bmatrix}.$ (2)
Second, a stable contact should guarantee that the movement of the gripper is synchronized with the handle’s motion. To establish a benchmark for stable contact $\mathcal{C}_{1}$, an initial gripper pose $\boldsymbol{T}_{1}$ is manually configured. This methodology is commonly employed in the manipulation of articulated objects [9]. A contact $\mathcal{C}_{i}$ is considered stable if its deviation from $\mathcal{C}_{1}$ stays within a specified threshold $d$:
$f_{d}(\mathcal{C}_{i},\mathcal{C}_{1})\leq d,$ (3)
where $f_{d}(\mathcal{C}_{i},\mathcal{C}_{1})$ computes the difference between $\mathcal{C}_{i}$ and $\mathcal{C}_{1}$.
@@
Upon establishing stable contact, the manipulation process commences with a preliminary direction relative to frame $\{i\}$ for operating the articulated object. This direction, represented as $r$, is encapsulated by a rotation matrix $\boldsymbol{R}^{i}\in SO(3)$. A detailed discussion on obtaining it and its impact on the efficiency of the method is provided in Section VI-B. Then, the homogeneous transformation matrix $\boldsymbol{T}_{r}^{i}$, resulting from moving along $r$ by a distance $t$, is defined as:
$\boldsymbol{T}_{r}^{i}=\begin{bmatrix}\boldsymbol{I}_{3\times 3}&\boldsymbol{R}^{i}\boldsymbol{t}\\ \begin{matrix}0&0&0\end{matrix}&1\end{bmatrix},$ (4)
where $\boldsymbol{t}$ denotes the translation vector whose Euclidean norm is equal to $t$:
$\left\|\boldsymbol{t}\right\|_{2}=t.$ (5)
The robot follows the preliminary direction, executing a series of gripper frame poses $\boldsymbol{T}_{1:n}$ to guide the handle’s transition from $J_{s}$ to $J_{e}$ while ensuring continuous stable contact. For each incremental step $i=1,\dots,n-1$, the next pose $\boldsymbol{T}_{i+1}$ is computed as follows:
$\boldsymbol{T}_{i+1}&=\boldsymbol{T}_{i}\boldsymbol{T}_{i+1}^{i}\\ &=\boldsymbol{T}_{i}\underbracket{\boldsymbol{T}_{r}^{i}}_{\text{execution}}\underbracket{\operatorname*{arg\,min}_{\boldsymbol{T}^{r}_{i+1}\in SE(3)}\left\|f_{c}(\boldsymbol{T}_{i}\boldsymbol{T}_{r}^{i}\boldsymbol{T}^{r}_{i+1},J)-\mathcal{C}_{1}\right\|_{2}}_{\text{recovery}}\end{aligned}.$ (6)
@@
The execution term in Eq. 6 involves proceeding to move along the preliminary direction as described in Eq. 4. This navigation progresses under the assumption of no prior knowledge about the object handle’s trajectory, posing a risk of exceeding the tolerance limits necessary for maintaining stable contact. To balance this risk and maintain efficiency, it is crucial to carefully select the magnitude of $t$ in Eq. 4. The objective is to maximize $t$ while ensuring that the constraints for stable contact are met. This optimization can be mathematically represented as:
@@
subject to $f_{e}(f_{c}(\boldsymbol{T}_{i}\boldsymbol{T}_{r}^{i},J),\mathcal{C}_{0})\leq\boldsymbol{e},$
$f_{s}(f_{c}(\boldsymbol{T}_{i}\boldsymbol{T}_{r}^{i},J),\mathcal{C}_{0})\leq s,$
$f_{d}(f_{c}(\boldsymbol{T}_{i}\boldsymbol{T}_{r}^{i},J),\mathcal{C}_{1})\leq d.$
@@
The recovery term in Eq. 6 implies that at each step, the robot adjusts its pose to minimize any deviation from the stable contact state $\mathcal{C}_{1}$, which may arise due to the execution step.